OpenStudy (anonymous):

Consider two functions f (x) = x^2 and g (x) = Sqrt(x) find fog (x) ?

OpenStudy (anonymous):

f o g (x) means f(g(x))..substitute the value of your g(x) to the x of f(x) make sense?

OpenStudy (anonymous):

\[\sqrt{x}\] is it the right answer?

OpenStudy (anonymous):

no... \(\sqrt x\) is g(x). since f(x) = x^2...substitute g(x) to x^2 f o g(x0 = f(g(x)) = g(x)^2 = (sqrt x)^2 you getting it?

OpenStudy (anonymous):

yes I got it, thank you.